ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 257, 2023

Bayer-Fluckiger, E. and T.-Y. Lee Norm tori of étale algebras and unramified Brauer groups

1

Bodin, A. and P. Dèbes Coprime values of polynomials in several variables

25

Breuer, T. and R. M. Guralnick Finite groups can be generated by a π-subgroup and a π′-subgroup

55

Cadoret, A., C. Y. Hui and A. Tamagawa ℚ_{ℓ}- versus 𝔽_{ℓ}- coefficients in the Grothendieck–Serre/Tate conjectures

71

Centeleghe, T. G. and J. Stix Categories of abelian varieties over finite fields II: Abelian varieties over 𝔽_{q} and Morita equivalence

103

Chatzidakis, Z. and N. Ramsey Measures on perfect e-free PAC fields

171

Chemotti, F., J. Mináč, T. T. Nguyen, A. Schultz, J. Swallow and N. D. Tân Quaternion algebras and square power classes over biquadratic extensions

193

Dèbes, P. See A. Bodin

Efrat, I. The kernel generating condition and absolute Galois groups  217

Esnault, H. and M. Kerz Local systems with quasi-unipotent monodromy at infinity are dense

251

Fried, S. and D. Haran Θ-Hilbertianity and strong Θ-Hilbertianity  263

Gajda, W. and S. Petersen Local to global principles for homomorphisms of abelian schemes

281

Guralnick, R. M. See T. Breuer

Haran, D. and P. A. Zalesskii Relatively projective pro-p groups  313

Haran, D. See S. Fried Harbater, D., J. Hartmann and D. Krashen Bounding cohomol-

ogy classes over semiglobal fields

353

Hartmann, J. See D. Harbater Heller, L., J. Mináč, T. T. Nguyen, A. Schultz and N. D. Tân

Galois module structure of some elementary p-abelian extensions

389

Hui, C. Y. See A. Cadoret Im, B.-H., M. Larsen and S. Zhan A hyperelliptic curve mapping

to specified elliptic curves

409

König, J. and D. Neftin The Hilbert–Grunwald specialization property over number fields 433 Kerz, M. See H. Esnault Krashen, D. See D. Harbater Larsen, M. See B.-H. Im Lee, T.-Y. See E. Bayer-Fluckiger

Lubotzky, A. and D. Neftin Sylow-conjugate number fields  465

Malliaris, M. and S. Shelah Shearing in some simple rank one theories

481

Mináč, J. See F. Chemotti Mináč, J. See L. Heller

Moret-Bailly, L. A henselian preparation theorem  519

Neftin, D. See A. Lubotzky Neftin, D. See J. König Nguyen, T. T. See F. Chemotti Nguyen, T. T. See L. Heller Paolini, G. and S. Shelah On the existence of uncountable Hopfian

and co-Hopfian abelian groups

533

Petersen, S. See W. Gajda Pop, F. On the Pythagoras number of function fields of curves over

number fields

561

Ramsey, N. See Z. Chatzidakis Razon, A. Primitive recursive decidability for large rings of algebraic

integers inside the compositum of all symmetric extensions of ℚ

575

Schneps, L. The Fay relations satisfied by the elliptic associator  591

Schultz, A. See F. Chemotti Schultz, A. See L. Heller Shelah, S. See G. Paolini Shelah, S. See M. Malliaris Stix, J. See T. G. Centeleghe Swallow, J. See F. Chemotti Tamagawa, A. See A. Cadoret Tân N. D. See F. Chemotti Tân N. D. See L. Heller Zalesskii, P. A. See D. Haran Zhan, S. See B.-H. Im